Parkins Company produces and sells a single product. The companys income statement for the most recent month is given below: $240,000 Sales (6,000 units at $40 per unit) Less Manufacturing costs: Direct materials Direct labor (variable) Variable factory overhead Fixed factory overhead Gross Margin Less selling and other expenses Variable selling and other expenses 24,000 Fixed selling and other expenses 42,000 $48,000 60,000 12,000 30,000 150,000 90,000 Net operating income $24,000 Required: a. Compute the companys monthly break-even point in units of product units b. What would the companys net operating income be if sales increased by 25% and there is no change in fixed expenses? $ c. What dollar sales must the company achieve in order to earn a net operating income of $50,000 per month? $
Answer #1

a) Break even point in units = Fixed cost / Contribution margin per unit

Variable cost = 48000+60000+12000+24000 = 144000

Contribution margin per unit = (240000-1440000)/6000 = 16 per unit

Fixed cost = 30000+42000 = 72000

Break even point in units = 72000/16 = 4500 units

b) Net operating income if sales increase by 25%

Sales = 240000*1.25 = 300000

Contribution margin ratio = 144000/240000 = 60%

Contribution margin = 300000*60% = 180000

Net operating income = 180000-72000 = 108000

c) Required dollar sales = (72000+50000)/.60 = 203333